OpenStudy (anonymous):

Convert the mixed number to an improper fraction. 6 7/8

OpenStudy (anonymous):

8x6 = 48 + 7 = 55/8. I do believe I did that correctly.

OpenStudy (black_ninja123):

Yes that is correct

OpenStudy (anonymous):

thanks